GENERAL CHEMISTRY I Stoichiometry and Reactions NAME: ux 1.hat are the reactants in the chemical equation below ? Grade Agts) AuNO, (aq) ANO, (aq) Au(s) a, Silver &Gold Nitrate . Silver& Gold b. Gold&Silver Nitrate d. Gold Nitrate & Silver Nitrate 2. Report the molar mass of Auso 3. How many hydrogen molecules are contained in 4.0 grams of H ? a. o.4M b. 12.0x10 c. 4.Omol d. 2.0g 4 What is the mass of 1.00 mole of calcium fluoride, CaF,? a. 78.1g Circ 2Fe (s) 0.2 mol Circle the compound below that you would expect to be insoluble in water. a. MgsO, Balance the reaction shown below. b. 40.08 amu c. 19.0Omol d. 38.0 g/mol 5. le the substance that is the limiting reagent in the reaction below. 3Nİ (NO3) 2 (s) 0.2 mol 2Fe (NO,), (ag) 6. b PbCl, c. KS d. Ba (OH), 7. N,P2 (s) Tell whether or not you would expect each compound below to be an acid, neither acid nor base Mg (OH) 2 A compound has an empirical formula of C,N.H and a molecular mass of abou CsCi Circle the correct molecular formula of the compound b. CNgH I is an ionic solid; in aqueous solution the ions formed will be a. Co2 & I A compound consisting of only oxygen and sulfur has a mass percentag Col d. Co sulfur. Predict the correct empirical formula. d. so b. SO2 c. S202 carbon, C(s) and 3 mol hydrogen, H2(g). (HINT: this is a combination a. 1.Omol Oxygen and hydrogen react according to the equation, 2H2(g)+02(g) Predict the theoretical yield of methane, CH, (g), that can be produce d. 3.0 L b. 18g c. 258 ? mnl ef hydrogen are reacted with 1 mol of oxygen and 15 grams lno0 than, greater than or equal to
Answer #1

1. The reactant in the chemical reaction below

(a) Silver and AuNO3 (reactants are always written on the left side of the reaction arrow)

2. The molar mass of AUSO4 is

Molar Mass of Au = 79 Molar Mass of S = 32   Molar mass of O=16

∴ total mass of AuSO4 = 79 + 32 + 4× 16 = 175g

3. mass of 1 mole of H2 is 2g which contain 6.023 × 1023 molecules . therefore 4 gm of H2 will have 2 moles therefore number of Hydrogen molecule will be 2 × 6.023 × 1023 =      12.046 × 1023

4. mass of 1 mole of CaF2 is equal to molar mass of CaF2 .

Molar mass of CaF2 = 40.1 + 19×2 = 78.1 g